### v3

# FedRef: Communication-Efficient Bayesian Fine-Tuning using a Reference Model

Abstract
Federated learning (FL) collaboratively trains artificial intelligence (AI) models to ensure user data privacy. Sharing only model updates generated from local training on client data with the server enhances user data privacy. However, model performance may suffer due to data and system heterogeneity among clients in FL scenarios. Previous studies have proposed model optimization, fine-tuning, and personalization to achieve improved model performance. Despite these efforts, models resulting from FL scenarios often exhibit catastrophic forgetting, which increases the communication and computational costs of clients for model optimization and raises energy consumption. To address these challenges, we propose a reference model-based fine-tuning method for federated learning that overcomes catastrophic forgetting in each round. Our method is derived from Bayesian parameter-efficient transfer learning and includes an proximal term. It employs a reference model that incorporates previous model parameters and reviews previous global features in the model optimization step to mitigate catastrophic forgetting. As a result, our method achieves higher model performance and lower communication and computational costs for clients than existing methods.
I Introduction
Federated learning (FL) has recently been proposed as a promising solution to protect user data privacy while allowing collaborative model training between independent clients (devices or institutions) [yang2019federated][lim2020federated]. User data are fundamentally protected because clients share only updated local model parameters (resulting from local training using local user data) and other non-private values with the server or other clients. The server then computes the global model by aggregating the local model updates received from the clients. However, as highlighted in [fedavg], FL approaches face numerous challenges, including:
  1. 1.
    Optimization of predictive model performance.
  2. 2.
    Decreasing computational costs and energy consumption.
  3. 3.
    Protecting the global model from malicious users and clients.
Our focus is on optimizing predictive model performance and reducing computational costs on client devices.
In the related work section, we discuss FedAvg, FedProx, and FedOpt, which are existing approaches for model optimization in FL scenarios, along with their limitations regarding client-side computational costs and catastrophic forgetting [mccloskey1989catastrophic][french1999catastrophic] [goodfellow2013empirical]. We also explain Bayesian Transfer Learning. We then present FedRef, a novel, communication-efficient Bayesian fine-tuning approach that leverages a reference model. FedRef mitigates catastrophic forgetting in each round by integrating features from previous models while performing maximum a posteriori (MAP) estimation.
In earlier work, MAP estimation has been used to optimize an objective function for transfer learning between a target model with pre-trained parameters and a source model [bayesian]. In contrast, traditional model optimization in FL scenarios employs a proximal term to minimize the deviation of a local model from the global model [fedprox][reddi2021adaptivefederatedoptimization]. This proximal term can also be considered when performing MAP estimation. In our research, we use MAP estimation to optimize an objective function that is updated using aggregated model information from previous rounds exclusively on the server side, thereby minimizing the computational costs for clients.
In our experiments, we consider two types of tasks: multi-class image classification and semantic segmentation on medical images. For multi-class image classification, we use two data sets: FEMNIST and CINIC-10. For semantic segmentation on medical images, we employ the FeTS2022 dataset, which is partitioned under a distributed, non-IID (Non-Independent and Identically Distributed) setting. In the semantic segmentation task, each client represents a hospital, as shown in Fig. 1.
For the segmentation task, we present the predictive model performance across different rounds using graphical representations, based on evaluation metrics commonly used in medical image segmentation studies [karimi2019reducing, chen2021transunet, hatamizadeh2022unetr, zhu2022medical]: mean Intersection over Union (mIoU), Dice Coefficient, and Hausdorff Distance.
In summary, the proposed FedRef method demonstrates high predictive performance on both medical and general computer vision tasks while minimizing computational costs for clients.
Figure 1: Basic medical FedRef system.
II Related Work
II-1 FedAvg
The authors of [fedavg] propose FedAvg, a typical federated learning system designed for communication-efficient training of deep neural networks on decentralized data. FedAvg demonstrates higher predictive performance than FedSGD on the MNIST classification task under both IID and non-IID conditions using the Federated Averaging method. Federated Averaging is based on stochastic gradient descent (SGD) [chen2016revisiting] and aggregates updates from $K$ client models by weighting them according to the number of samples per client, as shown in equation (1).
$\theta_{r+1}=\sum_{k=1}^{K}\frac{n_{k}}{n}\theta_{k}^{r}$ (1)
$r$: communication round, $\theta_{k}$ : model parameters of client $k$, $n_{k}$: number of samples of client $k$, $n$: total number of client samples. We also adopted this type of aggregation for our FedRef method; however, other weighted aggregation strategies can be easily integrated.
On the client side, the local training process is defined by Equation (2).
$\theta_{i}^{(s+1)}=\theta_{k}^{(s)}-\eta\nabla F_{k}(\theta_{k}^{(s)};\mathcal{B}_{k}^{(s)})$ (2)
$s$ : local training step, $\eta$: learning rate, $\mathcal{B}_{k}^{(s)}$: mini-batch sampled from the local dataset of client $k$ in step $s$, $F_{k}(\theta_{k}^{(s)};\mathcal{B}_{k}^{(s)})$: local loss function evaluated on a mini-batch.
Algorithm 1 illustrates the complete FedAvg process flow.
Algorithm 1 FedAvg: Communication-Efficient Learning of Deep Networks from Decentralized Data
1:Initialize global model weights $\theta_{0}$
2:for each round $r=1,2,\dots,R$ do
3:  Server selects clients $\mathcal{S}_{r}\subseteq\{1,\dots,K\}$
4:  for all client $k\in\mathcal{S}_{r}$ in parallel do
5:   Client $k$ receives $\theta_{r}$ from the server
6:   Client $k$ performs local training:
7:    $\theta_{r+1}^{k}\leftarrow\text{ClientUpdate}(k,\theta_{r})$
8:  end for
9:  Server aggregates the client updates:
10:   $\theta_{r+1}\leftarrow\sum_{k\in\mathcal{S}_{r}}\frac{n_{k}}{n_{\mathcal{S}_{r}}}\theta_{r+1}^{k}$
11:   $n_{\mathcal{S}_{r}}=\sum_{k\in\mathcal{S}_{r}}n_{k}$
12:end for
II-2 FedProx
The authors of [fedprox] study federated optimization in heterogeneous networks. In the local objective function that is optimized on each client, they introduce a “Proximal Term”, as shown in Equation (3), ensuring that local models do not deviate significantly from the global model. This approach enables model optimization that accounts for data skew and heterogeneity. However, it increases the computational costs of clients, as clients must compute the proximal term and send it to the server.
$\min_{\theta_{k}}\left(\mathbf{E}_{\mathcal{D}_{k}}\left[\mathcal{L}(\theta_{k},\mathcal{D}_{k})\right]+\frac{\mu}{2}\|\theta_{k}-\theta_{\text{global}}\|^{2}\right)$ (3)
$k$ : client index, $\theta_{k}$ : local model parameters of client $k$, $\mathcal{L}(\theta_{k},\mathcal{D}_{k})$ : cost function evaluated on the local data $\mathcal{D}_{k}$ of client $k$, $\theta_{\text{global}}$ : global model parameters, $\mu$: hyperparameter that represents the proximal term strength, $\|\theta_{k}-\theta_{\text{global}}\|^{2}$: the L2 (Euclidean) distance between the local model parameters and the global model parameters.
Algorithm 2 illustrates the approach followed by FedProx.
Algorithm 2 FedProx: Federated Optimization in Heterogeneous Networks
1:Initialize global model weights $\theta_{0}$
2:for each round $r=1,2,\dots,R$ do
3:  Server selects clients $\mathcal{S}_{r}\subseteq\{1,\dots,K\}$
4:  for all client $k\in\mathcal{S}_{r}$ in parallel do
5:   Client $k$ receives $\theta_{r}$ from the server
6:   Client $k$ performs local training:
7:    $\theta_{r+1}^{k}\leftarrow\text{ClientUpdate\_Prox}(k,\theta_{r},\mu)$
8:  end for
9:  Server aggregates the client updates:
10:   $\theta_{r+1}\leftarrow\sum_{k\in\mathcal{S}_{r}}\frac{n_{k}}{n_{\mathcal{S}_{r}}}\theta_{r+1}^{k}$
11:end for
On the server side, model parameters are only aggregated and returned to the clients. This leads to the issue of concentrated computational costs on the client side. For further details, on the client side, Algorithm 3 introduces local training with a proximal term, which further increases the computational costs for clients.
Algorithm 3 ClientUpdate_Prox($k$, $\theta_{r}$, $\mu$)
1:Compute cost and proximal term:
[…]
3:Update local model:
[…]
5:return $\theta$
II-3 FedOpt
The authors of [reddi2021adaptivefederatedoptimization] focus on integrating adaptive optimizers such as Adam, Yogi, and Adagrad at the server side, leading to variants of FedAvg that are known as FedAdam, FedYogi, and FedAdagrad, respectively. Following the FedAdagrad equation (4):
$g_{r}=\frac{1}{K}\sum_{k=1}^{K}\Delta\theta_{r}^{(k)},$
$G_{r}=G_{r-1}+g_{r}^{2},$
$\theta_{r+1}=\theta_{r}-\eta\cdot\frac{g_{r}}{\sqrt{G_{r}}+\tau}$ (4)
$g_{r}$: averaged update from all participating clients, $G_{r}$: accumulated sum of squared gradients (maintained on the server), $\tau$: small constant for numerical stability (e.g., $10^{-6}$). Following the FedAdam equation (II-3):
$m_{r}=\beta_{1}m_{r-1}+(1-\beta_{1})g_{r},$
$v_{r}=\beta_{2}v_{r-1}+(1-\beta_{2})g_{r}^{2},$
$\hat{m}_{r}=\frac{m_{r}}{1-\beta_{1}^{r}},\,\hat{v}_{r}=\frac{v_{r}}{1-\beta_{2}^{r}},$
$\theta_{r+1}=\theta_{r}-\eta\cdot\frac{\hat{m}_{r}}{\sqrt{\hat{v}_{r}}+\tau}$ (5)
$m_{r}$: first moment vector (mean of gradients), $v_{r}$: second moment vector (uncentered variance of gradients), $\hat{m}_{r},\hat{v}_{r}$: bias-corrected moment estimates, $\beta_{1},\beta_{2}$: decay rates for first and second moments (e.g., $\beta_{1}=0.9,\beta_{2}=0.999$).
Finally, following the FedYogi equation (II-3):
$v_{r}=v_{r-1}-(1-\beta_{2})\cdot sign(v_{r-1}-g_{r}^{2})\cdot g_{r}^{2},$
$m_{r}=\beta_{1}m_{r-1}+(1-\beta_{1})g_{r},$
$\hat{m}_{r}=\frac{m_{r}}{1-\beta_{1}^{r}}\,\hat{v}_{r}=\frac{v_{r}}{1-\beta_{2}^{r}},$
$w_{r+1}=w_{r}-\eta\cdot\frac{\hat{m}_{r}}{\sqrt{\hat{v}_{r}}+\tau}$ (6)
sign()($\cdot$): Element-wise sign function.
For details, we present the complete FedOpt approach in Algorithm 4:
Algorithm 4 FedOpt: Adaptive Federated Optimization
1:Input: $\theta_{0}$, CLIENTOPT, SERVEROPT
2:
3:for $r=0,\dots,R-1$ do
4:  Sample a subset $\mathcal{S}$ of clients
5:  $\theta_{r,0}^{i}=\theta_{r}$
6:  for each client $i\in\mathcal{S}$ in parallel do
7:   for $k=0,\dots,K-1$ do
8:     Compute $g_{r,k}^{i}$ of $\nabla F_{i}(\theta_{r,k}^{i})$
9:     $\theta_{r,k+1}^{i}=\texttt{CLIENTOPT}(\theta_{r,k}^{i},g_{r,k}^{i},\eta_{l},r)$
10:   end for
11:   $\Delta_{r}^{i}=\theta_{r,K}^{i}-\theta_{r}$
12:  end for
13:  $\Delta_{r}=\frac{1}{|\mathcal{S}|}\sum_{i\in\mathcal{S}}\Delta_{r}^{i}$
14:  $\theta_{r+1}=\texttt{SERVEROPT}(\theta_{r},-\Delta_{r},\eta,r)$
15:end for
FedAdam, FedYogi, and FedAdagrad also require client-side computations to optimize their proximal term. Among similar state-of-the-art strategies, recent research such as FedDyn [jin2023feddyn] has been proposed, which adaptively selects the server-side optimizer. However, both FedOpt and FedDyn still incur client-side computational costs due to the use of a proximal term.
II-4 Bayesian Transfer Learning
The authors of [bayesian] present a Bayesian transfer learning approach for mitigating catastrophic forgetting and formulate the optimization of the model fine-tuning cost as performing Maximum A Posteriori (MAP) estimation, as shown in Equation(7):
$\theta^{*}=\arg\max_{\theta}\left[\log p(D_{B}|\theta)+\log p(\theta|D_{A})\right]$ (7)
$D_{A}$: data used for pretraining, $D_{B}$: data used for downstream fine-tuning.
In Equation (7), $\log p(\theta|D_{A})$ can be approximated by Laplace Approximation [mackay1992practical] to yield Equation (8):
$\log p(\theta|D_{A})\approx f(\theta_{0})+\tfrac{1}{2}(\theta-\theta_{0})^{\top}H(\theta-\theta_{0})$ (8)
$H$: negative Hessian of the log-posterior, approximated by the Fisher Information Matrix (FIM).
Finally, the authors of [bayesian] make use of Equation (9) to define the objective function for the optimization of transfer learning:
$\mathcal{L}(\theta)=\underbrace{L_{B}(\theta)}_{\text{task loss}}+\lambda\,(\theta-\theta_{0})^{\top}F\,(\theta-\theta_{0})$ (9)
$\lambda$: regularization strength, $F$: FIM. To decrease computational costs and simplify the FIM to an identity matrix, Equation (9) can be simplified to Equation (10).
$\mathcal{L}_{\text{l2-sp}}(\theta)=L_{B}(\theta)+\lambda\|\theta_{k}-\theta_{0,k}\|^{2}$ (10)
Despite being simplified, Equation (10) has been shown to increase robustness against catastrophic forgetting in transfer learning [xuhong2018explicit][kirkpatrick2017overcoming]. Under alternative assumptions, Kronecker-factored approximation can also be applied [martens2015optimizing] [george2018fast], which accounts for the covariance between parameters within each layer.
As a result, this paper investigates the optimization of transfer learning in an FL scenario. Whereas conventional transfer learning happens between only two models – one a pre-trained model intended for transfer, and the other a target mode – in FL scenarios, the server must aggregate and optimize many different client models. Therefore, we need a new MAP estimation problem formulation for FL scenarios.
III FedRef: Communication-Efficient Bayesian Fine Tuning using a Reference Model
For optimal model performance and low client-side computational cost, we propose FedRef, a communication-efficient Bayesian fine-tuning method with a reference model that overcomes catastrophic forgetting by inferring from global models of previous rounds.
Figure 2: A reference model for inferring with alternative global models.
The reference model, in Fig. 2, perform as a standard(target) model of proximal term objective reference to mitigate a deviation from previous round models. In the existing study [bayesian], transferred learning is a concept of model distillation and focus on how do distillate to transfer features from alternative reference models to a target model. The most of FL scenario also focus on how do distillate to transfer features from each clients to global model. Therefore, in the most of FL scenarios, the ideal situation is to the client’s model inference as an independent event and when the model inference of all clients is, on average, the most accurate phase.
Building on the Bayesian fine-tuning approach presented in [bayesian], evaluating the proximal term of the global model is formulated as performing MAP estimation in the context of FL scenarios, as shown in Equation (11):
$\theta^{*}=\arg\max_{\theta}[\log p(\theta|D_{ref})+\log p(D_{1}|\theta)+$
$logp(D_{2}|\theta)+\dots+\log p(D_{K}|\theta)]$ (11)
Selected client numbers are denoted as $k\in[1,2,3,..,K]$, where $K$ is the total number of selected clients. $D_{ref}$: represents data for definition involved a reference model.
$D_{ref}$ is defined solely to explain our optimal MAP problem in Equation (III). This approach helps overcome catastrophic forgetting in each round by integrating features from previous rounds into the MAP problem and subsequently optimizing the integrated MAP value.
Finally, the objective function derived in Equation (11) can be expressed by combining Equations (8, 9, and 10) to form Equation (III).
$L_{ref}(\theta^{1},\theta^{2},\forall F)=\sum_{k}^{K}F_{k}\cdot\mathrm{diag}(W_{1},\dots,W_{K})$
$+\lambda(\theta^{1}_{r}-\theta^{1}_{r-1})^{2}+\lambda(\theta^{2}_{r}-\theta^{2}_{r-1})^{2}$ (12)
In Equation (III), the constant term $\sum_{k}^{K}F_{k}$ denotes the sum of client losses. The diagonal matrix $\mathrm{diag}(W_{1},\dots,W_{K})$ represents the aggregation weights (e.g., $\frac{n_{k}}{n}$). The $\theta^{1}$ is a global model and $\theta^{2}$ is a reference model. The term $(\theta_{r}-\theta_{r-1})^{2}$ signifies the $L_{2}$ regularization of the model parameters $\theta_{r}-\theta_{r-1}$.
Regarding parameter requirements, the proposed FedRef method only needs client losses. Model optimization is solely performed on the server side, which helps reduce the client-side computational cost, because clients only perform local training. In FedRef, $\theta^{2}$ should be set as the reference model $\theta_{ref}$. For details on $\theta_{ref}$, the reference model is defined as follows:
$\theta_{r+1}^{2}=\text{A}(\theta^{1}_{r-p},\theta^{1}_{r-p+1},\dots,\theta^{1}_{r})$ (13)
where $p$ is the number of selected subsets from previous global models, which can be set heuristically, and the function A, representing global features from the previous rounds, is calculated as
$\text{A}(\forall\theta)=\sum_{i=1}^{p}\frac{1}{p}\theta^{1}_{r+1-i}$ (14)
A suitable value for $p$ was determined heuristically to be between 3 and 5. However, the choice should also take into account the available memory resources for storing model parameters.
We describe FedRef in detail in Algorithm 5:
Algorithm 5 FedRef: Communication-Efficient Bayesian Fine-Tuning with a Reference Model
1:Client:
2:for epoch do
3:  for batch do
4:   $\theta_{k}^{(s+1)}=\theta_{k}^{(s)}-\eta\nabla F_{k}(\theta_{k}^{(s)};\mathcal{B}_{k}^{(s)})$
5:  end for
6:end for
7:send $\theta$ , $F(\theta^{(s)};\mathcal{B}^{(s)})$ to Server
8:
9:Server:
10:for round do
11:$K$-client select
[…]
15: broadcast $\theta$ to Clients
16:end for
17:
18:Aggregation($\forall\theta)$:
[…]
20:return $\theta$
21:
22:Bayesian Optimization($\theta^{1},\theta^{2},\forall F$):
[…]
24:return $\theta^{1}$
25:
[…]
IV Experiments
IV-A Experimental Setup
IV-A1 Global
We developed a federated learning system using the Flower framework [beutel2020flower]. The detailed experimental conditions are summarized in Table I. Figure 3 illustrates the data partitioning and the number of samples assigned to each client in the FL setting.
TABLE I: Experimental Setup
Setting Specification
FL Framework Flower: a friendly federated learning framework
Language Python
Operating System Linux 24.04 LTS
GPU NVIDIA RTX 4090
Tools Visual Studio Code
Within Python, the Flower framework enables the easy construction of federated learning servers and clients from high-level code. It supports both Object-Oriented Programming (OOP) and Functional Programming (FP) paradigms for privacy-preserving machine learning strategies (e.g., FedAvg, FedProx, and FedOpt). The source code for our method is publicly available at https://github.com/TaehwanY98/Fed-Ref.
In the classification tasks, the objective function is the asymmetric loss [ridnik2021asymmetric], whereas in the segmentation task, the objective function is the focal dice loss [lin2017focal] [yeung2022unified], used to mitigate class imbalance.
Figure 3: Number of samples per client for each dataset.
IV-A2 FEMNIST
The FEMNIST dataset is an image classification benchmark consisting of handwritten digits (0-9), lowercase letters (a-z), and uppercase letters (A-Z). This results in a total of 62 unique labels. Each sample in the dataset includes a $\times28$ grayscale image along with a writer-id, hsf-id, and character label. The writer-id is a unique identifier for each person who wrote the characters, which allows for the creation of a non-independent and identically distributed (non-IID) data split – a crucial aspect of realistic federated learning simulations. The dataset is available through the Hugging Face Hub, making it easy to integrate with the Flower federated learning framework[DBLP:journals/corr/abs-1812-01097][DBLP:journals/corr/abs-2007-14390]. For our experiments, we partitioned the data according to the hsf-id information.
IV-A3 FeTs2022
The FeTs2022 dataset refers to the dataset used in the Federated Tumor Segmentation (FeTS) Challenge 2022, which focused on brain tumor segmentation using federated learning approaches. The dataset contains multimodal brain MRI scans of glioma patients. Each patient folder typically includes the following four MRI modalities: T1-weighted, T1-weighted post-contrast, T2-weighted, and T2 Fluid-Attenuated Inversion Recovery (FLAIR). The data labels provide segmentations for three tumor sub-regions: enhancing tumor, tumor core, and whole tumor. The dataset can be downloaded from the official Synapse site [fets2022].
To evaluate FedRef, we perform the segmentation task on the FeTs2022 dataset, which features a specific non-IID data partitioning, as shown in Fig. 3. The baseline model used for this task is a 3D U-Net.
IV-A4 CINIC-10
The CINIC-10 dataset is a large-scale image classification dataset specifically designed to serve as a drop-in replacement for the CIFAR-10 dataset in federated learning research. It was created to address the need for a more challenging benchmark dataset that fills the gap between the relatively small CIFAR-10 dataset and the much larger ImageNet dataset. The CINIC-10 dataset contains a total of 270,000 images, which is 4.5 times more than CIFAR-10. The images are all $\times32$ pixels, making them directly compatible with models designed for CIFAR-10 [DBLP:journals/corr/abs-2007-14390][darlow2018cinic10imagenetcifar10].
Table II compares the different types of client-side computations and the communicated resources for each method. FedRef clients perform local training and communicate model parameters and single cost values to the server. The combination of low client-side computational costs and low communication requirements enables efficient communication with fast local training.
TABLE II: Comparison of client computations and communicated resources per method
Method Client computations Communicated resources Server computations
FedAvg Local training Model param Aggregation
FedProx Local training, Proximal term Model param Aggregation
FedOpt Local training, Proximal term Model param, Proximal term Aggregation, Model optimization
FedRef Local training Model param, Cost value Aggregation, Model Bayesian optimization
TABLE III: FEMNIST classification: required rounds
Method $AsymmetricLoss<0.2$
FedAvg 21 rounds
FedProx 20 rounds
FedOpt 21 rounds
FedRef 20 rounds
TABLE IV: CINIC-10 classification: required rounds
Method $AsymmetricLoss<0.4$ $F1score>18\%$
FedAvg 28 rounds 22 rounds
FedProx 28 rounds 22 rounds
FedOpt 28 rounds 22 rounds
FedRef 23 rounds 23 rounds
TABLE V: FeTS2022 segmentation: required rounds
Method $mIOU>14\%$ $Hausdorff95<9$
FedAvg 30 rounds 23 rounds
FedProx 27 rounds 18 rounds
FedOpt 28 rounds 19 rounds
FedRef 27 rounds 17 rounds
In Tables III and IV, we applied a tiny convolutional neural network (tiny-CNN) as the baseline model for small-sized images to minimize overfitting on the FEMNIST and CINIC-10 datasets. As a result, our FedRef approach achieves a training speed comparable to other optimization strategies (FedProx, FedOpt) under the tiny-CNN setting, despite the reduced client-side computational costs.
Furthermore, in Table V, FedRef achieves better model performance, with a higher mean intersection of union (mIoU) and a lower Hausdorff distance than other methods on medical images exhibiting data heterogeneity and data skew. Since the cost function corresponds to the Dice score relative to the evaluation metric (mIoU), FedRef effectively optimizes and minimizes the cost value. Consequently, the performance of the FedRef model is strongly influenced by the cost function.
Figure 4: Graph of server-side centralized evaluation on each strategy.
V Challenges
FedRef still faces several challenges, including the selection of hyperparameters (such as the client-server learning rate and the constant value representing the total number of selected models from previous rounds) and ensuring protection against malicious users and model inference attacks. As different assumptions can be made regarding the FIM, we plan to investigate various FIM formulations in future work to identify a more suitable proximal term.
VI Conclusions
For optimizing model training in FL scenarios, existing studies have proposed various proximal terms that serve as regularization terms in the objective function, preventing local models from drifting too far from the global model. However, these approaches do not address the issue of catastrophic forgetting in each round, and their proximal terms have to be computed on the client side, increasing client-side computational costs. Our FedRef method, a communication-efficient Bayesian fine-tuning approach using a reference model, overcomes these limitations through server-side optimization of an objective function that includes a proximal term inferring previous global features. This approach leads to low client-side computational cost and high model optimization performance, thereby improving model training efficiency in federated learning settings.