### v4

# FedRef: Communication-Efficient Bayesian Fine-Tuning using a Reference Model

Abstract
Federated learning (FL) collaboratively trains artificial intelligence (AI) models to ensure user data privacy. Sharing only model updates generated from local training on client data with the server enhances user data privacy. However, model performance may suffer due to data and system heterogeneity among clients in FL scenarios. Previous studies have proposed model optimization, fine-tuning, and personalization to achieve improved model performance. Despite these efforts, models resulting from FL scenarios often exhibit catastrophic forgetting, which increases the communication and computational costs of clients for model optimization and raises energy consumption. To address these challenges, we propose a reference model-based fine-tuning method for federated learning that overcomes catastrophic forgetting in each round. Our method is derived from Bayesian parameter-efficient transfer learning and includes an proximal term. It employs a reference model that incorporates previous model parameters and reviews previous global features in the model optimization step to mitigate catastrophic forgetting. As a result, our method achieves higher model performance and lower communication and computational costs for clients than existing methods.
I Introduction
Federated learning (FL) has recently been proposed as a promising solution to protect user data privacy while allowing collaborative model training between independent clients (devices or institutions) [yang2019federated][lim2020federated]. User data are fundamentally protected because clients share only updated local model parameters (resulting from local training using local user data) and other non-private values with the server or other clients. The server then computes the global model by aggregating the local model updates received from the clients. However, as highlighted in [fedavg], FL approaches face numerous challenges, including:
  1. 1.
    Optimization of predictive model performance.
  2. 2.
    Decreasing computational costs and energy consumption.
  3. 3.
    Protecting the global model from malicious users and clients.
Our focus is on optimizing predictive model performance and reducing computational costs on client devices.
Critical issues and Bayesian optimization. In the existing studies section, we discuss FedAvg, FedProx, and FedOpt, which are existing approaches for model optimization in FL scenarios, along with their limitations regarding client-side computational costs and catastrophic forgetting [mccloskey1989catastrophic][french1999catastrophic] [goodfellow2014empirical]. We also explain Bayesian Transfer Learning. We then present FedRef, a novel, communication-efficient Bayesian fine-tuning approach that leverages a reference model. FedRef mitigates catastrophic forgetting in each round by integrating features from previous models while performing maximum a posteriori (MAP) estimation.
A definition of MAP estimation In earlier work, MAP estimation has been used to optimize an objective function for transfer learning between a target model with pre-trained parameters and a source model [bayesian]. In contrast, traditional model optimization in FL scenarios employs a proximal term to minimize the deviation of a local model from the global model [fedprox]. This proximal term can also be considered when performing MAP estimation. In our research, we use MAP estimation to optimize an objective function that is updated using aggregated model information from previous rounds exclusively on the server side, thereby minimizing the computational costs for clients.
Fig. 1 provides an overview of our proposed FedRef algorithm. The FedRef algorithm can be summarized in the following steps:
  1. 1.
    Server sends a model to clients.
  2. 2.
    Clients perform local training.
  3. 3.
    Clients send a model parameter and cost value.
  4. 4.
    Server aggregates selected clients parameters.
  5. 5.
    Server computes a reference model.
  6. 6.
    Server optimize global model derived from our Bayesian optimization using the reference model.
  7. 7.
    Server send the result of the optimization to the clients.
In our experiments, we consider two types of tasks: multi-class image classification and semantic segmentation on medical images. For multi-class image classification, we use two data sets: FEMNIST and CINIC-10. For semantic segmentation on medical images, we employ the FeTS2022 dataset, which is partitioned under a distributed, non-IID (Non-Independent and Identically Distributed) setting. In the semantic segmentation task, each client represents a hospital, as shown in Fig. 1.
For the segmentation task, we present the predictive model performance across different rounds using graphical representations, based on evaluation metrics commonly used in medical image segmentation studies [karimi2019reducing, chen2021transunet, hatamizadeh2022unetr, zhu2022medical]: mean Intersection over Union (mIoU), Dice Coefficient (DC), and Hausdorff Distance . And we present required rounds on reaching specific evaluation scores by Tables for each datasets and methods. And then finally, we demonstrate communication efficiency from our FedRef method by charts including relative communication resource ratios using required rounds on reaching specific evaluation scores.
In summary, the proposed FedRef:communication efficient Bayesian fine tuning with reference model method demonstrates high predictive performance on both medical and general computer vision tasks while minimizing computational costs for clients. Due to minimized computational costs and parameter-efficient fine-tuning (PEFT), We achieve and contribute higher communication efficiency on federated learning scenarios than existing studies.
Figure 1: Overview of the our FedRef: communication-efficient bayesian fine tuning using a reference model Algorithm.
II Existing studies
Federated Learning has emerged as a distributed machine learning paradigm enabling collaborative model training without centralized data collection, thus preserving user privacy. Meanwhile, parameter-efficient fine-tuning (PEFT) techniques allow efficient adaptation of large-scale pre-trained models using only a fraction of their parameters. The combination of FL and PEFT represents a promising direction for privacy-preserving and resource-effective learning across distributed devices.
For the federated learning backgrounds and paradigms, we introduce basic federated learning papers about optimizing model parameters and fine tuning.
II-1 FedAvg
The authors of [fedavg] propose FedAvg, a typical federated learning system designed for communication-efficient training of deep neural networks on decentralized data. FedAvg demonstrates higher predictive performance than FedSGD on the MNIST classification task under both IID and non-IID conditions using the Federated Averaging method. Federated Averaging is based on stochastic gradient descent (SGD) [chen2016revisiting] and aggregates updates from $K$ client models by weighting them according to the number of samples per client, as shown in Equation (1).
$\theta_{r+1}=\sum_{k=1}^{K}\frac{n_{k}}{n}\theta_{k}^{r}$ (1)
$r$: communication round, $\theta_{k}$ : model parameters of client $k$, $n_{k}$: number of samples of client $k$, $n$: total number of client samples. We also adopted this type of aggregation for our FedRef method; however, other weighted aggregation strategies can be easily integrated.
On the client side, the local training process is defined by Equation (2).
$\theta_{i}^{(s+1)}=\theta_{k}^{(s)}-\eta\nabla F_{k}(\theta_{k}^{(s)};\mathcal{B}_{k}^{(s)})$ (2)
$s$ : local training step, $\eta$: learning rate, $\mathcal{B}_{k}^{(s)}$: mini-batch sampled from the local dataset of client $k$ in step $s$, $F_{k}(\theta_{k}^{(s)};\mathcal{B}_{k}^{(s)})$: local loss function evaluated on a mini-batch.
Algorithm 1 illustrates the complete FedAvg process flow.
Algorithm 1 FedAvg: Communication-Efficient Learning of Deep Networks from Decentralized Data
1:Initialize global model weights $\theta_{0}$
2:for each round $r=1,2,\dots,R$ do
3:  Server selects clients $\mathcal{S}_{r}\subseteq\{1,\dots,K\}$
4:  for all client $k\in\mathcal{S}_{r}$ in parallel do
5:   Client $k$ receives $\theta_{r}$ from the server
6:   Client $k$ performs local training:
7:    $\theta_{r+1}^{k}\leftarrow\text{ClientUpdate}(k,\theta_{r})$
8:  end for
9:  Server aggregates the client updates:
10:   $\theta_{r+1}\leftarrow\sum_{k\in\mathcal{S}_{r}}\frac{n_{k}}{n_{\mathcal{S}_{r}}}\theta_{r+1}^{k}$
11:   $n_{\mathcal{S}_{r}}=\sum_{k\in\mathcal{S}_{r}}n_{k}$
12:end for
II-2 FedProx
The authors of [fedprox] study federated optimization in heterogeneous networks. In the local objective function that is optimized on each client, they introduce a “Proximal Term”, as shown in Equation (3), ensuring that local models do not deviate significantly from the global model. This approach enables model optimization that accounts for data skew and heterogeneity. However, it increases the computational costs of clients, as clients must compute the proximal term and send it to the server.
$\min_{\theta_{k}}\left(\mathbf{E}_{\mathcal{D}_{k}}\left[\mathcal{L}(\theta_{k},\mathcal{D}_{k})\right]+\frac{\mu}{2}\|\theta_{k}-\theta_{\text{global}}\|^{2}\right)$ (3)
$k$ : client index, $\theta_{k}$ : local model parameters of client $k$, $\mathcal{L}(\theta_{k},\mathcal{D}_{k})$ : cost function evaluated on the local data $\mathcal{D}_{k}$ of client $k$, $\theta_{\text{global}}$ : global model parameters, $\mu$: hyperparameter that represents the proximal term strength, $\|\theta_{k}-\theta_{\text{global}}\|^{2}$: the L2 (Euclidean) distance between the local model parameters and the global model parameters.
Algorithm 2 illustrates the approach followed by FedProx.
Algorithm 2 FedProx: Federated Optimization in Heterogeneous Networks
1:Initialize global model weights $\theta_{0}$
2:for each round $r=1,2,\dots,R$ do
3:  Server selects clients $\mathcal{S}_{r}\subseteq\{1,\dots,K\}$
4:  for all client $k\in\mathcal{S}_{r}$ in parallel do
5:   Client $k$ receives $\theta_{r}$ from the server
6:   Client $k$ performs local training:
7:    $\theta_{r+1}^{k}\leftarrow\text{ClientUpdate\_Prox}(k,\theta_{r},\mu)$
8:  end for
9:  Server aggregates the client updates:
10:   $\theta_{r+1}\leftarrow\sum_{k\in\mathcal{S}_{r}}\frac{n_{k}}{n_{\mathcal{S}_{r}}}\theta_{r+1}^{k}$
11:end for
On the server side, model parameters are only aggregated and returned to the clients. This leads to the issue of concentrated computational costs on the client side. For further details, on the client side, Algorithm 3 introduces local training with a proximal term, which further increases the computational costs for clients.
Algorithm 3 ClientUpdate_Prox($k$, $\theta_{r}$, $\mu$)
1:Compute cost and proximal term:
[…]
3:Update local model:
[…]
5:return $\theta$
II-3 FedOpt
The authors of [reddi2020adaptive] focus on integrating adaptive optimizers such as Adam, Yogi, and Adagrad at the server side, leading to variants of FedAvg that are known as FedAdam, FedYogi, and FedAdagrad, respectively. Following the FedAdagrad equation (4):
$g_{r}=\frac{1}{K}\sum_{k=1}^{K}\Delta\theta_{r}^{(k)},$
$G_{r}=G_{r-1}+g_{r}^{2},$
$\theta_{r+1}=\theta_{r}-\eta\cdot\frac{g_{r}}{\sqrt{G_{r}}+\tau}$ (4)
$g_{r}$: averaged update from all participating clients, $G_{r}$: accumulated sum of squared gradients (maintained on the server), $\tau$: small constant for numerical stability (e.g., $10^{-6}$).
Following the FedAdam Equation (II-3):
$m_{r}=\beta_{1}m_{r-1}+(1-\beta_{1})g_{r},$
$v_{r}=\beta_{2}v_{r-1}+(1-\beta_{2})g_{r}^{2},$
$\hat{m}_{r}=\frac{m_{r}}{1-\beta_{1}^{r}},\,\hat{v}_{r}=\frac{v_{r}}{1-\beta_{2}^{r}},$
$\theta_{r+1}=\theta_{r}-\eta\cdot\frac{\hat{m}_{r}}{\sqrt{\hat{v}_{r}}+\tau}$ (5)
$m_{r}$: first moment vector (mean of gradients), $v_{r}$: second moment vector (uncentered variance of gradients), $\hat{m}_{r},\hat{v}_{r}$: bias-corrected moment estimates, $\beta_{1},\beta_{2}$: decay rates for first and second moments (e.g., $\beta_{1}=0.9,\beta_{2}=0.999$).
Finally, following the FedYogi Equation (II-3):
$v_{r}=v_{r-1}-(1-\beta_{2})\cdot sign(v_{r-1}-g_{r}^{2})\cdot g_{r}^{2},$
$m_{r}=\beta_{1}m_{r-1}+(1-\beta_{1})g_{r},$
$\hat{m}_{r}=\frac{m_{r}}{1-\beta_{1}^{r}}\,\hat{v}_{r}=\frac{v_{r}}{1-\beta_{2}^{r}},$
$w_{r+1}=w_{r}-\eta\cdot\frac{\hat{m}_{r}}{\sqrt{\hat{v}_{r}}+\tau}$ (6)
sign()($\cdot$): Element-wise sign function.
For details, we present the complete FedOpt approach in Algorithm 4:
Algorithm 4 FedOpt: Adaptive Federated Optimization
1:Input: $\theta_{0}$, CLIENTOPT, SERVEROPT
2:
3:for $r=0,\dots,R-1$ do
4:  Sample a subset $\mathcal{S}$ of clients
5:  $\theta_{r,0}^{i}=\theta_{r}$
6:  for each client $i\in\mathcal{S}$ in parallel do
7:   for $k=0,\dots,K-1$ do
8:     Compute $g_{r,k}^{i}$ of $\nabla F_{i}(\theta_{r,k}^{i})$
9:     $\theta_{r,k+1}^{i}=\texttt{CLIENTOPT}(\theta_{r,k}^{i},g_{r,k}^{i},\eta_{l},r)$
10:   end for
11:   $\Delta_{r}^{i}=\theta_{r,K}^{i}-\theta_{r}$
12:  end for
13:  $\Delta_{r}=\frac{1}{|\mathcal{S}|}\sum_{i\in\mathcal{S}}\Delta_{r}^{i}$
14:  $\theta_{r+1}=\texttt{SERVEROPT}(\theta_{r},-\Delta_{r},\eta,r)$
15:end for
FedAdam, FedYogi, and FedAdagrad also require client-side computations to optimize their proximal term. Among similar state-of-the-art strategies, recent research such as FedDyn [jin2023feddyn] has been proposed, which adaptively selects the server-side optimizer. However, both FedOpt and FedDyn still incur client-side computational costs due to the use of a proximal term.
II-4 Bayesian Transfer Learning
The authors of [bayesian] present a Bayesian transfer learning approach for mitigating catastrophic forgetting and formulate the optimization of the model fine-tuning cost as performing Maximum A Posteriori (MAP) estimation, as shown in Equation(7):
$\theta^{*}=\arg\max_{\theta}\left[\log p(D_{B}|\theta)+\log p(\theta|D_{A})\right]$ (7)
$D_{A}$: data used for pretraining, $D_{B}$: data used for downstream fine-tuning.
In Equation (7), $\log p(\theta|D_{A})$ can be approximated by Laplace Approximation [mackay1992practical] to yield Equation (8):
$\log p(\theta|D_{A})\approx f(\theta_{0})+\tfrac{1}{2}(\theta-\theta_{0})^{\top}H(\theta-\theta_{0})$ (8)
$H$: negative Hessian of the log-posterior, approximated by the Fisher Information Matrix (FIM).
Finally, the authors of [bayesian] make use of Equation (9) to define the objective function for the optimization of transfer learning:
$\mathcal{L}(\theta)=\underbrace{L_{B}(\theta)}_{\text{task loss}}+\lambda\,(\theta-\theta_{0})^{\top}F\,(\theta-\theta_{0})$ (9)
$\lambda$: regularization strength, $F$: FIM. To decrease computational costs and simplify the FIM to an identity matrix, Equation (9) can be simplified to Equation (10).
$\mathcal{L}_{\text{l2-sp}}(\theta)=L_{B}(\theta)+\lambda\|\theta_{k}-\theta_{0,k}\|^{2}$ (10)
Despite being simplified, Equation (10) has been shown to increase robustness against catastrophic forgetting in transfer learning [xuhong2018explicit][kirkpatrick2017overcoming]. Under alternative assumptions, Kronecker-factored approximation can also be applied [martens2015optimizing] [george2018fast], which accounts for the covariance between parameters within each layer.
As a result, this paper investigates the optimization of transfer learning in an FL scenario. Whereas conventional transfer learning happens between only two models – one a pre-trained model intended for transfer, and the other a target mode – in FL scenarios, the server must aggregate and optimize many different client models. Therefore, we need a new MAP estimation problem formulation for FL scenarios. For completeness, the mathematical background on Laplacian regularization in Bayesian optimization and a comparativediscussion with FedDyn are provided in the Appendix.
III FedRef: Communication-Efficient Bayesian Fine Tuning using a Reference Model
The existing studies have limitations of high client-side computational costs and catastrophic forgetting issues on each rounds. For optimal model performance and low client-side computational cost, we propose FedRef, a communication-efficient Bayesian fine-tuning method with a reference model that overcomes catastrophic forgetting by inferring from global models of previous rounds.
Figure 2: A reference model for inferring with alternative global models.
III-A Descriptions of a reference model
As illustrated in Fig. 2, the reference model serves as a target model of proximal term objective reference to mitigate a deviation from previous round models. In the existing study [bayesian], transfer learning is a concept of model distillation and focus on how to distill and transfer features from alternative reference models to a target model. The most of FL scenario also focus on how do distillate to transfer features from each clients to global model. Therefore, in the most of FL scenarios, the ideal case assumes each client’s model inference is an independent event and when the model inference of all clients is, on average, the most accurate phase.
III-B A MAP estimation in FL scenarios
Building on the Bayesian fine-tuning approach presented in [bayesian], evaluating the proximal term of the global model is formulated as performing MAP estimation in the context of FL scenarios, as shown in Equation (11):
$\theta^{*}=\arg\max_{\theta}[\log p(\theta|D_{ref})+\log p(D_{1}|\theta)+$
$logp(D_{2}|\theta)+\dots+\log p(D_{K}|\theta)]$ (11)
Selected client numbers are denoted as $k\in[1,2,3,..,K]$, where $K$ is the total number of selected clients. $D_{ref}$: represents data for definition involved a reference model.
$D_{ref}$ is defined solely to explain our optimal MAP problem in Equation (III-C). This approach helps overcome catastrophic forgetting in each round by integrating features from previous rounds into the MAP problem and subsequently optimizing the integrated MAP value.
III-C A objective estimation derived from the MAP estimation
Equation (11) reformulates global optimization in an FL setting as a MAP estimation, where the prior term $\log\,p(\theta|D_{ref})$ incorporates historical knowledge from previous global models, and the likelihood terms $\log\,p(D_{k}|\theta)$ accumulate evidence from the currently participating clients. Finally, the objective function derived in Equation (11) can be expressed by combining Equations (8, 10, and 12) to form Equation (III-C). Likelihood approximation via local loss is denoted at Equation 12,13. Equation (11) expresses the MAP objective in a federated setting, where the server combines a prior derived from a reference model with the likelihood contributions from all participating clients. Since the server cannot access raw client data, each client’s negative log-likelihood is approximated by its local loss function.
$-\log p(D_{k}\mid\theta)\approx F_{k}(\theta)$ (12)
Equation (12) formalizes this approximation, allowing the likelihood term to be replaced with a computable local loss.
$\sum_{k=1}^{K}\log p(D_{k}\mid\theta)\Longleftrightarrow-\sum_{k=1}^{K}F_{k}(\theta)$ (13)
Based on this, Equation (13) shows that the global likelihood can be written as a weighted sum of client losses, where the weights reflect the relative data sizes of each client. This produces a practical formulation that aligns with standard FL aggregation.
To incorporate knowledge from previous rounds, FedRef introduces a reference model that serves as the prior mean. Using a Laplace approximation, the prior becomes an $L_{2}$ regularization term encouraging the global model to remain close to this reference. This helps stabilize optimization and mitigates catastrophic forgetting in non-IID environments.
Finally, Equation (III-C) results from combining the likelihood approximation and the prior regularization. This objective balances client empirical loss with temporal consistency, enabling FedRef to perform stable and communication-efficient global updates.
$L_{ref}(\theta^{1},\theta^{2},\forall F)=\sum_{k}^{K}F_{k}\cdot\mathrm{diag}(W_{1},\dots,W_{K})$
$+\lambda(\theta^{1}_{r}-\theta^{1}_{r-1})^{2}+\lambda(\theta^{2}_{r}-\theta^{2}_{r-1})^{2}$ (14)
In Equation (III-C), $F_{k}$: local loss value from client $k$, the constant term $\sum_{k}^{K}F_{k}$ denotes the sum of client losses. The diagonal matrix $\mathrm{diag}(W_{1},\dots,W_{K})$ represents the aggregation weights (e.g., $\frac{n_{k}}{n}$). The $\theta^{1}$ is a global model and $\theta^{2}$ is a reference model. The term $(\theta_{r}-\theta_{r-1})^{2}$ signifies the $L_{2}$ regularization of the model parameters $\theta_{r}-\theta_{r-1}$.
III-D Formal definition of a reference model
Regarding parameter requirements, the proposed FedRef method only needs client losses. Model optimization is solely performed on the server side, which helps reduce the client-side computational cost, because clients only perform local training. In FedRef, $\theta^{2}$ should be set as the reference model $\theta_{ref}$. For details on $\theta_{ref}$, the reference model is defined as follows:
$\theta_{r+1}^{2}=\text{A}(\theta^{1}_{r-p},\theta^{1}_{r-p+1},\dots,\theta^{1}_{r})$ (15)
where $p$ is the number of selected subsets from previous global models, which can be set heuristically, and the function A, representing global features from the previous rounds, is calculated as
$\text{A}(\forall\theta)=\sum_{i=1}^{p}\frac{1}{p}\theta^{1}_{r+1-i}$ (16)
A suitable value for $p$ was determined heuristically to be between 3 and 5. However, the choice should also take into account the available memory resources for storing model parameters.
The resulting objective $L_{ref}$ thus combines the weighted sum of client-side losses with a temporally regularized penalty that constrains the drift of global parameters, effectively balancing adaptation and retention across communication rounds.
The server computes the gradient of $L_{ref}$ and performs a single Bayesian optimization step, as detailed in Algorithm 5 (lines 17–18). This step updates the global parameters $\theta^{1}$ by jointly considering aggregated client losses ($F_{k}$) and reference constraints ($\theta^{2}$).
Algorithm 5 FedRef: Communication-Efficient Bayesian Fine-Tuning with a Reference Model
1:Client:
2:for epoch do
3:  for batch do
4:   $\theta_{k}^{(s+1)}=\theta_{k}^{(s)}-\eta\nabla F_{k}(\theta_{k}^{(s)};\mathcal{B}_{k}^{(s)})$
5:  end for
6:end for
7:send $\theta$ , $F(\theta^{(s)};\mathcal{B}^{(s)})$ to Server
8:
9:Server:
10:for round do
11:$K$-client select
[…]
15: broadcast $\theta$ to Clients
16:end for
17:
18:Aggregation($\forall\theta)$:
[…]
20:return $\theta$
21:
22:Bayesian Optimization($\theta^{1},\theta^{2},\forall F$):
[…]
24:return $\theta^{1}$
25:
[…]
IV Experiments Analysis
IV-A Experimental Setup
IV-A1 Environments
We developed a federated learning system using the Flower framework [beutel2020flower]. The detailed experimental conditions are summarized in Table I. Figure 3 illustrates the data partitioning and the number of samples assigned to each client in the FL setting.
TABLE I: Experimental Setup
Setting Specification
FL Framework Flower: a friendly federated learning framework
Language Python
Operating System Linux 24.04 LTS
GPU NVIDIA RTX 4090
Tools Visual Studio Code
Within Python, the Flower framework enables the easy construction of federated learning servers and clients from high-level code. It supports both Object-Oriented Programming (OOP) and Functional Programming (FP) paradigms for privacy-preserving machine learning strategies (e.g., FedAvg, FedProx, and FedOpt). The source code for our method is publicly available at MyGithub.
In the classification tasks, the objective function is the asymmetric loss [ridnik2021asymmetric], whereas in the segmentation task, the objective function is the focal dice loss [lin2017focal] [yeung2022unified], used to mitigate class imbalance.
Figure 3: Number of samples per client for each dataset.
IV-A2 FEMNIST
The FEMNIST dataset is an image classification benchmark consisting of handwritten digits (0-9), lowercase letters (a-z), and uppercase letters (A-Z). This results in a total of 62 unique labels. Each sample in the dataset includes a $\times28$ grayscale image along with a writer-id, hsf-id, and character label. The writer-id is a unique identifier for each person who wrote the characters, which allows for the creation of a non-independent and identically distributed (non-IID) data split – a crucial aspect of realistic federated learning simulations. The dataset is available through the Hugging Face Hub, making it easy to integrate with the Flower federated learning framework[DBLP:journals/corr/abs-1812-01097][DBLP:journals/corr/abs-2007-14390]. For our experiments, we partitioned the data according to the hsf-id information.
IV-A3 CINIC-10
The CINIC-10 dataset is a large-scale image classification dataset specifically designed to serve as a drop-in replacement for the CIFAR-10 dataset in federated learning research. It was created to address the need for a more challenging benchmark dataset that fills the gap between the relatively small CIFAR-10 dataset and the much larger ImageNet dataset. The CINIC-10 dataset contains a total of 270,000 images, which is 4.5 times more than CIFAR-10. The images are all $\times32$ pixels, making them directly compatible with models designed for CIFAR-10 [DBLP:journals/corr/abs-2007-14390][darlow2018cinic10imagenetcifar10].
IV-A4 FeTs2022
The FeTs2022 dataset refers to the dataset used in the Federated Tumor Segmentation (FeTS) Challenge 2022, which focused on brain tumor segmentation using federated learning approaches. The dataset contains multimodal brain MRI scans of glioma patients. Each patient folder typically includes the following four MRI modalities: T1-weighted, T1-weighted post-contrast, T2-weighted, and T2 Fluid-Attenuated Inversion Recovery (FLAIR). The data labels provide segmentations for three tumor sub-regions: enhancing tumor, tumor core, and whole tumor. The dataset can be downloaded from the official Synapse site [fets2022].
To evaluate FedRef, we perform the segmentation task on the FeTs2022 dataset, which features a specific non-IID data partitioning, as shown in Fig. 3. The baseline model used for this task is a 3D U-Net [cciccek20163d]. 3D U‑Net provides a powerful framework for volumetric medical image segmentation, extending convolutional feature extraction into three dimensions to exploit full 3D context. Despite its computational demands, it has become a foundation for modern medical imaging pipelines and continues to inspire improved architectures for efficient and precise 3D segmentation.
IV-B Experimental Results
Evaluations of predictive model performances. Table II compares the different types of client-side computations and the communicated resources and server-side computations for each method. FedRef clients perform local training and communicate model parameters and cost values to the server. The combination of low client-side computational costs and low communication requirements enables efficient communication with fast local training.
TABLE II: Comparison of client computations and communicated resources per method
Method Client computations Communicated resources Server computations
FedAvg Local training Model param Aggregation
FedProx Local training, Proximal term Model param Aggregation
FedOpt Local training, Proximal term Model param, Proximal term Aggregation, Model optimization
FedRef Local training Model param, Cost value Aggregation, Model Bayesian optimization
An ablation study of FedRef. We conduct an ablation study on the two main hyperparameters of FedRef: the regularization strength $\lambda$ and the number of reference models $p$. As shown in Table III, $\lambda$ controls the influence of the reference model during optimization. Larger $\lambda$ values strengthen temporal regularization, and $\lambda$ between 0.0001 and 0.001 provides stable convergence across all datasets. For FEMNIST and CINIC‑10, $\lambda$ = 0.001 yields the best performance, while smaller $\lambda$ values perform slightly better on FeTS2022 due to stronger non‑IID conditions.
The parameter $p$ determines how many past global models are averaged to build the reference model. While larger $p$ increases temporal stability, it also raises memory usage. In practice, $p$ = 3 offers a good balance between performance and computational cost, and higher values do not yield additional improvements.
These results confirm that both $\lambda$ and $p$ play complementary roles: $\lambda$ governs the strength of reference-based regularization, and $p$ controls the temporal depth of the reference model. Proper tuning of these parameters enables FedRef to maintain stable optimization while preserving communication and computational efficiency.
TABLE III: FedRef heuristic parameters analysis: total round 30
$\lambda$ $p$ FEMNIST: AsymmLoss CINIC-10: AsymmLoss FeTs2022: DC FeTs2022: HF95
0.0001 3 13.95 9.245 14.09% 8.685
0.0001 5
0.0005 3 13.95 8.860 14.09% 8.676
0.0005 5
0.001 3 14.09 8.392 13.98% 8.847
0.001 5
Evaluations of communication efficiency. This section evaluates how efficiently FedRef reduces the communication cost required to reach a given performance level, compared to baseline optimization strategies such as FedAvg, FedProx, and FedOpt. Because FedRef does not require clients to compute or transmit proximal terms, the amount of data communicated per round remains similar to FedAvg. However, by performing Bayesian optimization with a reference model entirely on the server side, FedRef accelerates convergence and therefore lowers the total number of communication rounds needed. Fig. 4. reports the relative communication resource ratio, computed using Equation (17), for the FEMNIST, CINIC-10, and FeTS2022 benchmarks. This ratio incorporates both the number of bytes communicated per round ($b$) and the required rounds ($\gamma$) to reach predefined evaluation thresholds. A lower value indicates higher communication efficiency.
As summarized in Tables IVVI, FedRef consistently requires fewer rounds to reach the target performance across all benchmarks. On FEMNIST, FedRef requires 12 rounds to reach an Asymmetric Loss <14.5<14.5, outperforming FedAvg (13 rounds) and FedProx (16 rounds). On CINIC-10, FedRef achieves the Asymmetric Loss <9.5<9.5 threshold in 17 rounds, the best among all methods (1–3 rounds faster than others). On FeTS2022, FedRef reaches a Dice-Coefficient $14\%$ in 24 rounds, outperforming FedAvg (27 rounds) and achieving the target where FedOpt and FedProx fail.
TABLE IV: FEMNIST classification: required rounds
Method $AsymmetricLoss<14.5$ $Accuracy>9\%$
FedAvg 13 rounds 12 rounds
FedProx 16 rounds no meet the expectation
FedOpt 11 rounds 12 rounds
FedRef 12 rounds 12 rounds
TABLE V: CINIC-10 classification: required rounds
Method $AsymmetricLoss<9.5$ $Accuracy>10\%$
FedAvg 20 rounds 22 rounds
FedProx 19 rounds 22 rounds
FedOpt 18 rounds 22 rounds
FedRef 17 rounds 22 rounds
TABLE VI: FeTS2022 segmentation: required rounds
Method Dice-Coefficient $14\%$ Hausdorff95 <9<9
FedAvg 27 rounds 19 rounds
FedProx no meet the expectation 28 rounds
FedOpt no meet the expectation 20 rounds
FedRef 24 rounds 16 rounds
An evaluation metric definition. In the Fig. 4, relative communication resource ratios on each datasets are computed derived from equation (17):
$E_{max}(b,\gamma,\epsilon)=\text{max}(2\times b\times\gamma)$
$E_{min}(b,\gamma,\epsilon)=\text{min}(2\times b\times\gamma)$
$E=\frac{E_{s}-E_{min}}{E_{max}-E_{min}}+\epsilon$ (17)
The $E$ is a list of relative communication resource ratios on each federated learning strategies. $b$ includes the bytes of communication resource parameters (model parameter,… etc) on each federated learning strategies. $\gamma$ is a required rounds of specific standard of evaluation. $\epsilon$ can be set as a constant value for visualization in the Fig. 4. As a result, our proposed FedRef achieves higher communication efficient on reaching required rounds than existing works derived from the Fig. 4.
For further detail, in the Fig. 4, on FEMNIST, required round $\gamma$ represents $\gamma\in[11,12,13,16]$ derived from the Table IV if we define the standard of evaluation score to asymmetric loss. And $p$ includes 8.2 mega bytes from the global model. For the $\epsilon$ setting, we set $\epsilon$ to 1 in our experiments. In the Fig. 4, on CINIC-10 also, required round $\gamma$ represents $\gamma\in[17,18,19,20]$ derived from the Table V. And $p$ includes 8.3 mega bytes from the global model. For the $\epsilon$, we set $\epsilon$ to 1 in our experiments. In the Fig. 4, on FeTs2022 lastly, required round $\gamma$ represents $\gamma\in[16,19,20,28]$ derived from the Table V if we define the standard of evaluation score to mean Hausdorff distances. And $p$ includes 8.3 mega bytes from the global model. For the $\epsilon$, we set $\epsilon$ to 1 in our experiments.
Summary of results. In Tables IV and V, we applied a tiny convolutional neural network (tiny-CNN) as the baseline model for small-sized images to minimize overfitting on the FEMNIST and CINIC-10 datasets. As a result, our FedRef approach achieves a training speed comparable to other optimization strategies (FedProx, FedOpt) under the tiny-CNN setting, despite the reduced client-side computational costs. Furthermore, in Table VI, FedRef achieves better model performance, with both higher mean intersection of union (mIoU) and mean dice-coefficient (DC) and a lower Hausdorff distance than other methods on medical images exhibiting data heterogeneity and data skew. Since the cost function corresponds to the Dice score relative to the evaluation metric (DC), FedRef effectively optimizes and minimizes the cost value. Therefore, the performance of the FedRef model is strongly influenced by the cost function.
Consequently, FedRef reduces the total communication burden without increasing client-side computation, offering a practical and scalable optimization strategy for real-world federated learning scenarios.
Figure 4: Relative communication resource ratios on each datasets
Figure 5: Graph of server-side centralized evaluation on each strategy.
V Discussion and Future Work
FedRef still faces several challenges, including the selection of hyperparameters (such as the client-server learning rate and the constant value representing the total number of selected models from previous rounds) and ensuring protection against malicious users and model inference attacks. As different assumptions can be made regarding the FIM, we plan to investigate various FIM formulations in future work to identify a more suitable proximal term. Moreover, for preventing unbounded drift issues from non-IID federated learning scenarios which can be described as a scenario where this divergence is not naturally limited or controlled, potentially leading to performance degradation, slow convergence, or even model collapse, to mitigate the issues, alternative regularization techniques can be a solutions of the unbounded drift issues. Moreover, while FedRef reduces the client-side burden, it implicitly assumes that the reference model can be reconstructed from a fixed number of previous rounds ($p$). In highly dynamic client participation scenarios, this assumption may not hold, potentially affecting the stability of the Bayesian objective. Future studies could therefore explore adaptive reference model selection or hierarchical priors to further stabilize optimization without sacrificing efficiency.
VI Related work
The parameter efficient fine tuning (PEFT) studies contributed to decrease required rounds achieving specific evaluation scores and to minimize various non-IID issues where is occurred from real world federated learning scenarios. In the existing PEFT on federated learning studies [fedprompt], model freezing and compression methods can increase communication efficiency. And alternative optimization studies [li2018hyperband][jaderberg2017pbt][chen2025fedshaplex] also can be references of alternative objective estimations against our a MAP estimation. Furthermore, the existing personalized federated learning studies [scott2023pefll][zhang2025fedlth] also can be references for better predictive model performance on each client cases. In terms of communication efficient federated learning, the existing studies [sezgin2025energy][shi2025federated] not only consider minimizing computational costs and communication resources and compressing model, but also minimizing client and server side energy consumptions. For the studies [dou2025toward] [zhao2021sear] of protecting the global model from malicious users and clients, can be references of solution about filtering malicious users and clients and regularization of the malicious users and clients.
VII Conclusions
For optimizing model training in FL scenarios, existing studies have proposed various proximal terms that serve as regularization terms in the objective function, preventing local models from drifting too far from the global model. However, these approaches do not address the issue of catastrophic forgetting in each round, and their proximal terms have to be computed on the client side, increasing client-side computational costs. Our FedRef method, a communication-efficient Bayesian fine-tuning approach using a reference model, overcomes these limitations through server-side optimization of an objective function that includes a proximal term inferring previous global features. This approach leads to low client-side computational cost and high model optimization performance, thereby improving model training efficiency in federated learning settings.
Acknowledgment
This research was supported by the MSIT, Korea, NRF Korea (RS-2025-00557379, 50%) and the Innovative Human Resource Development for Local Intellectualization support program (IITP-2026-RS-2022-00156360, 30%) supervised by the IITP, and Convergence Security Core Talent Training Business Support Program (IITP-2026-RS-2024-00426853, 20%) supervised by the IITP.
Appendix A. Laplacian Expression in Bayesian Optimization
The Laplacian Expression in Bayesian Optimization: Bayesian Optimization (BO) seeks the global optimum of an unknown and expensive-to-evaluate objective function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ by constructing a probabilistic surrogate model and iteratively selecting evaluation points to balance exploration and exploitation. A common surrogate is the Gaussian Process (GP), which assumes that function values follow a joint Gaussian distribution characterized by a mean function $m(\mathbf{x})$ and a covariance (kernel) function $k(\mathbf{x},\mathbf{x}^{\prime})$.
Laplacian-Regularized Acquisition Optimization: To encourage spatially smooth exploration, the acquisition function $\alpha(\mathbf{x})$ can be regularized by the Laplacian to penalize irregular updates:
$\mathcal{L}(\alpha)=\int_{\Omega}\left\|\nabla\alpha(\mathbf{x})\right\|^{2}\,d\mathbf{x}=-\int_{\Omega}\alpha(\mathbf{x})\,\Delta\alpha(\mathbf{x})\,d\mathbf{x},$ (18)
where $\Omega$ denotes the search region. Minimizing $\mathcal{L}(\alpha)$ enforces smooth acquisition surfaces and stabilizes optimization over noisy, high–dimensional domains.
Laplace Operator in Bayesian Linear and Partial Differential Equation(PDE)-Constrained Models: In physics-informed or PDE-constrained Bayesian optimization, the Laplacian explicitly appears in the governing equation:
$\mathcal{A}(f)(\mathbf{x})=\Delta f(\mathbf{x})-g(\mathbf{x})=0,$ (19)
where $\mathcal{A}$ is the differential operator defining the system. Here, BO aims to infer or optimize $f$ under this partial differential constraint. The corresponding GP prior can be constructed by designing a covariance function $k$ that satisfies the same PDE constraint, i.e.,
$\mathcal{A}_{\mathbf{x}}\,\mathcal{A}_{\mathbf{x}^{\prime}}\,k(\mathbf{x},\mathbf{x}^{\prime})=\delta(\mathbf{x},\mathbf{x}^{\prime}),$ (20)
where $\delta(\cdot,\cdot)$ is the Dirac delta function.
Laplacian-based Transfer Regularization: To preserve smoothness and avoid negative transfer, some transfer learning and BO formulations impose a Laplacian regularization term.
$\Omega(f)=\frac{1}{2}\sum_{i,j}w_{ij}\bigl(f^{(i)}-f^{(j)}\bigr)^{2}=\mathbf{f}^{\top}L\mathbf{f},$ (21)
where L=DWL=D-W is the graph Laplacian derived from the inter-task similarity matrix $W$. This penalty enforces function consistency across related tasks and integrates naturally with GP priors or acquisition optimization.
Appendix B. FedDyn: Federated Learning with Dynamic Regularization
FedDyn: Federated Learning with Dynamic Regularization: (Federated Dynamics) is an optimization framework proposed to mitigate the challenges of data heterogeneity in Federated Learning (FL). In conventional FL approaches such as Federated Averaging (FedAvg) [fedavg], model divergence arises when local client updates differ significantly due to non-i.i.d. data distributions. FedDyn addresses this issue by introducing a dynamic regularization term that implicitly aligns local and global objectives, ensuring stable and fast convergence under client heterogeneity.
FedDyn modifies the local objective of each client to introduce a dynamic correction term, yielding the following local problem:
$\min_{\mathbf{w}_{k}}\bigl(f_{k}(\mathbf{w}_{k})-\boldsymbol{\lambda}_{k}^{\top}\mathbf{w}_{k}+\tfrac{\alpha}{2}\|\mathbf{w}_{k}\|^{2}\bigr),$ (22)
where $\alpha$ is a tunable regularization coefficient and $\boldsymbol{\lambda}_{k}$ is a dynamically updated gradient correction variable.
The global update is given by:
$\mathbf{w}^{t+1}=\frac{1}{K}\sum_{k=1}^{K}\bigl(\mathbf{w}_{k}^{t+1}-\tfrac{1}{\alpha}\boldsymbol{\lambda}_{k}^{\,t}\bigr),$ (23)
followed by
$\boldsymbol{\lambda}_{k}^{\,t+1}=\boldsymbol{\lambda}_{k}^{\,t}-\alpha(\mathbf{w}^{t+1}-\mathbf{w}_{k}^{t+1}).$ (24)
This coupling of local and global variables introduces a potential game structure that ensures consistency between global and local optimization dynamics.
The FedDyn formulation can be understood as constructing a penalized Lagrangian system between the global and local models:
$L(\{\mathbf{w}_{k}\},\mathbf{w},\{\boldsymbol{\lambda}_{k}\})=\sum_{k}p_{k}f_{k}(\mathbf{w}_{k})+\frac{\alpha}{2}\sum_{k}p_{k}\|\mathbf{w}_{k}-\mathbf{w}\|^{2}.$ (25)
The dynamic variable $\boldsymbol{\lambda}_{k}$ accumulates discrepancies between local and global models, thereby compensating for non‑i.i.d. data effects. When aggregated, these correction terms implicitly align all local objectives with the global one, leading to faster and more robust convergence than standard FedAvg or FedProx.